The measure of angle Y is twice the measure of its complement. Find the measure of complement of angle Y.: The measure of angle Y is twice the measure of its complement. Find the measure of complement of angle Y.
Answer by jim_thompson5910(9926)

Complementary angles add up to 90 degrees. So let x be the complement to angle y. Since "The measure of angle Y is twice the measure of its complement", this means that y=2x. Also, since the two angles add up to 90, this means that x+y=90


x+(2x)=90 Plug in y=2x into the second equation. In other words, replace each y with 2x. Notice we've eliminated the y variables. So we now have a simple equation with one unknown.


3x=90 Combine like terms on the left side


x=(90)/(3) Divide both sides by 3 to isolate x



x=30 Divide




Now that we know that x=30, we can plug this into y=2x to find y



y=2(30) Substitute 30 for each x


y=60 Simplify


So our answer is x=30 and y=60. So the measure of the complement of angle y is 30 degrees.
Question 147554: The measure of angle Y is twice the measure of its complement. Find the measure of complement of angle Y.: The measure of angle Y is twice the measure of its complement. Find the measure of complement of angle Y.
Answer by mangopeeler07(448)

complementary angles add up to 90 degrees. So first, set up and equation that says that The measure of angle Y plus its complement equals 90. Make sure that y represents y and not its complement. y+(1/2)y=90. Then combine like terms and get (3/2)y=90. Isolate the variable by multiplying both sides 2/3. You should get y=60. Again, that is y, and not its complement. So you can either do (1/2)y or (1/2)60 and get 30, or you can do 90-60 and get 30. Either way, thecomplementofy=30.